# Bounds on decaying sterile neutrinos via magnetic dipole moment from COB intensity

### v2

@@ Abstract
A recent observation by Long Range Reconnaissance Imager (LORRI) mounted on NASA’s New Horizons yielded the most accurate measurement of the cosmic optical background (COB). The reported COB intensity is $11.16\pm 1.35$ $\mathrm{nW/m^{2}/sr}$ at a pivot wavelength $\lambda_{piv}=0.608\,\mu\mathrm{m}$ observed in the range $0.4\,\mu\mathrm{m}\lesssim\lambda\lesssim 0.9\,\mu\mathrm{m}$. After subtracting the measured intensity from the deep Hubble Space Telescope count, diffused galactic light, and scattered light from bright star foregrounds, an anomalous intensity $2.99\pm 2.03\leavevmode\nobreak\ \mathrm{nW/m^{2}/sr}$ has been found. We considered radiatively decaying sterile neutrinos of mass scales keV and eV, as dark matter candidate and as dark radiation respectively, that could contribute to this anomalous reported intensity. Using this, we derive upper bounds on sterile-to-sterile and active-to-sterile transition magnetic dipole moments. We find that sterile neutrinos with mass of $\mathcal{O}(\rm keV)$ scale take values of the transition magnetic moment in the range $3\times 10^{-12}\,\mu_{B}-10^{-9}\,\mu_{B}$, whereas for mass of $\mathcal{O}(\rm eV)$ scale the transition magnetic moment in the range $3\times 10^{-13}\,\mu_{B}-10^{-12}\,\mu_{B}$, here $\mu_{B}$ is Bohr magneton.
Sterile-to-sterile transition magnetic dipole moment, Sterile-to-active transition magnetic dipole moment, Low Energy Effective Field Theory(LEFT)
@@ I Introduction
Line-intensity mapping (LIM) is an observational method for extragalactic astronomy and cosmology that measures the integrated intensity of spectral lines emitted from galaxies and the intergalactic medium at a specific observed frequency Kovetz et al. (2017, 2020). Since LIM experiments capture information from all incoming photons, they have the potential to detect electromagnetic radiation produced by dark matter decays directly. For instance, in Ref. Grin et al. (2007), authors have derived a stringent upper bound on the axion-photon coupling strength by searching for optical line emission produced from decaying axions. In Ref. Gong et al. (2016), authors have shown that photons produced from decaying axions with eV mass can explain the anisotropy of near-IR extragalactic background light. In Ref. Creque-Sarbinowski and Kamionkowski (2018), authors have proposed that LIM experiments can detect photons produced from decaying and annihilating dark matter particles. However, measuring the extragalactic background light in the optical band is a difficult task. The primary reason is mitigating the overwhelming foreground, which requires accurate modelling and well-calibrated instruments Bernal et al. (2022). In Ref. Zemcov et al. (2017a), authors have suggested observing such radiations in the optical spectrum—the cosmic optical background (COB) is possible if one can eliminate the foreground radiations, primarily solar in nature.
@@
The Long Range Reconnaissance Imager (LORRI) mounted on NASA New Horizons spacecraft, currently operating at a distance $51.3$ astronomical unit (AU) from the Sun, has been measuring the COB photon intensity in the wavelength range $0.4\,\mu\mathrm{m}\lesssim\lambda\lesssim 0.9\,\mu\mathrm{m}$ $(\sim 1.3\leavevmode\nobreak\ \rm eV-3\leavevmode\nobreak\ \,\rm eV)$ Zemcov et al. (2017b); Lauer et al. (2021, 2022); Symons et al. (2023); Postman et al. (2024). In $2021$, LORRI operating at 42-45 AU from the Sun reported an optical intensity from unknown sources of $8.8\pm 4.9\,\rm nW/m^{2}/sr$ $11.9\pm 4.6\,\rm nW/m^{2}/sr$ Lauer et al. (2021). Later, it yielded an anomalous intensity of $8.06\pm 1.92\,\rm nW/m^{2}/sr$ while operating at $51.3$ AU Lauer et al. (2022). In Lauer et al. (2022), the authors considered diffused galactic light (DGL) from the Milky Way and scattered light from bright stars (SSL) contributing to the total sky level. Further, they subtracted the intensity from integrated light of external galaxies (IGL), known as deep Hubble galaxy count, of $8.17\pm 1.18\,\rm nW/m^{2}/sr$ from the total COB level to report the anomalous intensity Lauer et al. (2022). Later, this result was also supported by an independent research following blind methodology Symons et al. (2023). However, in a recent analysis Postman et al. (2024), the authors have reworked the analysis of the foregrounds from the DGL and found the contribution of DGL to the COB intensity is greater compared to the one found in article Lauer et al. (2022). Thus yielding an anomalous intensity of $2.99\pm 2.03\,\rm{nW/m^{2}/sr}$ in the optical spectrum, relatively modest to the IGL intensity Postman et al. (2024).
@@
This anomalous photon excess may indicate the presence of physics beyond the Standard Model. Various studies have explored possible explanations, such as photons emitted from decaying axions with masses in the range of $\sim 5\leavevmode\nobreak\ \rm eV-25\leavevmode\nobreak\ \rm eV$ Bernal et al. (2022); Nakayama and Yin (2022); Carenza et al. (2023). Additionally, decaying sterile neutrinos could produce a similar photon signal through their decay into photons and neutrinos.
@@
Short-baseline neutrino experiments have hinted at the existence of additional neutrinos, and numerous studies have investigated sterile neutrinos as a potential dark matter candidate Ho and Scherrer (2013). The investigation of photon radiated from the sterile neutrinos can provide an exciting probe for the beyond standard model containing sterile neutrinos.
Motivated by these possibilities, we analyze photon flux originated from decaying sterile neutrinos. We also explore it as source of the observed anomalous intensity.
@@
In our work, we focus on the keV mass scale for sterile neutrinos, as they are considered a viable dark matter candidate, as well as we have considered the eV mass scale of decaying sterile neutrinos. Rather than considering a specific model, we adopt an effective field theory approach Beltrán et al. (2024), which makes our analysis more general and less dependent on particular model assumptions. Anomalies observed in short-baseline neutrino experiments, such as MiniBooNE Aguilar-Arevalo et al. (2021, 2018) and LSND Aguilar et al. (2001), have sparked significant interest in the possibility of eV-scale sterile neutrinos, which have been extensively studied in the literature. Motivated by these scenarios, we investigate the potential role of both the keV- and eV-scale sterile neutrinos in contributing to the reported anomalous COB intensity, thus deriving upper bounds on sterile-to-sterile and sterile-to-active transition magnetic dipole moment. Future experiments like the SPHEREx telescope surveying the sky at wavelengths between $0.75-5\,\mu m$ might shed light on the properties of radiatively decaying sterile neutrinos Crill et al. (2020); Doré et al. (2014, 2016, 2018).
@@
The paper is organised as follows. In Section II, we discuss the radiative decay process of sterile neutrinos through the sterile-to-sterile and sterile-to-active transition magnetic dipole moment in the low-energy effective field theory. In Section III, we derive the mean specific intensity for the keV mass range of decaying sterile neutrinos. In Section IV, we have analyzed our results by determining the minimal required magnetic dipole moment and decay width for sterile-to-sterile and active-to-sterile transition magnetic dipole moment and the existing limits on it. In this section, we have considered the two COB measurements. Here, we also derive the novel constraints on sterile-to-sterile transition magnetic dipole moment through the contribution on the COB from optical to X-rays region. In Section V, we present our conclusions with an outlook for future works.
@@
II Radiative Decay of Sterile Neutrino through Transition Magnetic Dipole Moment
@@
We consider Standard model augmented with sterile neutrinos in the effective field theory approach. Assuming sterile neutrino mass scales well below the electroweak scale, we can write the effective Lagrangian contributes to the magnetic moment of the neutrino sector:
[…]
@@
where, $d_{\rm NN\gamma}$ represents the sterile-to-sterile transition magnetic dipole moment and $d_{\rm\nu N\gamma}$ denotes the active-to-sterile transition magnetic dipole moment, respectively. The term $\mathcal{O}_{\rm NN\gamma}^{ij}=(\bar{N}_{Ri}^{c}\sigma_{\mu\nu}N_{Rj})F^{\mu\nu}$, $\mathcal{O}_{\nu N\gamma}^{\alpha i}=(\bar{\nu}_{L\alpha}^{c}\sigma_{\mu\nu}N_{Ri})F^{\mu\nu}$ represent the effective field operators, where $F^{\mu\nu}$ is the electromagnetic field strength tensor. The electric and magnetic field off-diagonal non-zero dipole moment components give rise to radiative decay through two different sterile states. Therefore, in this radiative decay process ($m_{1}\rightarrow m_{2}+\gamma$), the decay width of sterile neutrino can be written as Atre et al. (2009); Bondarenko et al. (2018); Balantekin and Vassh (2014)
@@
From Eq. (3), we can derive the sterile-to-sterile transition magnetic dipole moment $d_{\rm NN\gamma}$.
There are several experimental upper bounds on the neutrino magnetic moment. Recent studies of the NOMAND neutrino detector at CERN put constraints on the active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim(10^{-6}-10^{-9})\mu_{B}$ Gninenko and Krasnikov (1999). Here, $\mu_{B}\sim 3\times 10^{-7}\,\rm eV^{-1}$ represents Bohr magneton. In an anti-neutrino electron scattering experiment, by examining the electron recoil spectra, it was found that active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 10^{-9}\mu_{B}$ Derbin et al. (1993). The Super-Kamiokande experiment put an upper bound limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 1.1\times 10^{-10}\mu_{B}$ Liu et al. (2004), and the Borexino collaboration put the upper bound limit at $90\%$ confidence level as $5.4\times 10^{-11}\mu_{B}$ Agostini et al. (2017); Arpesella et al. (2008). MUNU collaboration has found the upper bound on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 9.0\times 10^{-11}\mu_{B}$ Babu et al. (2020), whereas TEXONO collaboration has obtained this active-to-sterile transition magnetic dipole moment upper bound limit $d_{\rm\nu N\gamma}\lesssim 7.4\times 10^{-11}\mu_{B}$ Balantekin and Vassh (2014). The Planck+BAO put the constraint on the active-to-sterile transition magnetic dipole moment, and it varies from $3.7\times 10^{-11}\mu_{B}$ to $9.1\times 10^{-12}\mu_{B}$ Carenza et al. (2024). Recent studies analysing the cooling of red giant stars have found an upper limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 1.2\times 10^{-12}\mu_{B}$ Capozzi and Raffelt (2020). The anomalous stellar cooling due to plasmon decay has been found an upper limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}$ is $d_{\rm\nu N\gamma}\lesssim 3\times 10^{-12}\mu_{B}$ Alexander (2016).
@@ III Mean specific intensity of decaying sterile neutrinos
where $h$ and $c$ are the Planck’s constant and the speed of light in vacuum, respectively. We define the second term on the right-hand side of Eq. (4) as $x^{-1}=\frac{2}{\left(1-m_{\xi}^{2}/m_{\chi}^{2}\right)}$. We note that for a two-photon decay process, $x=1/2$. The wavelength of photon produced from a particle decaying at redshift $z$ will be observed today as $\lambda_{obs}=(1+z)\lambda_{e}$. If these photons are not absorbed by the intergalactic gas along the light of sight, then they will be observed today as the extra-galactic background lights. The redshift dependence of the energy density of decaying dark matter particles is given by $\rho_{\chi}(t)=\rho_{\chi}(0)(1+z)^{3}\mathrm{e}^{-\Gamma_{\chi}t_{\rm U}}$, where $t_{\rm U}\sim 10^{18}\,\rm sec$ is the age of the universe, $\Gamma_{\chi}$ represents the decay rate, and $\rho_{\chi}(0)$ is the dark matter energy density today. If $\Gamma_{\chi}$ of decaying dark matter particle is such that $\Gamma_{\chi}t_{\rm U}\lesssim 10^{-2}$, we can then consider $e^{-\Gamma_{\chi}t_{\rm U}}\approx 1$. Later in this section, we have shown that the decay rate considered in this work justifies this approximation. The specific intensity of the photons $I_{\nu}$ observed at frequency $\nu_{\rm obs}$ today is given by Creque-Sarbinowski and Kamionkowski (2018)
[…]
where $\epsilon_{\nu_{\rm obs}}$ is the specific luminosity density and $H(z)$ is the Hubble parameter. For a Dirac delta decaying profile, $\epsilon_{\nu_{\rm obs}}$ is expressed as Chen and Kamionkowski (2004); Creque-Sarbinowski and Kamionkowski (2018)
@@
From the above equation, we can observe that $I_{\lambda}$ is directly proportional to the decay rate and the energy of the released photon. However, it is inversely related to the mass of the decaying particle and decays down over redshifts due to the expansion of the Universe.

### v3

# compat=1.1.0

@@ Abstract
A recent observation by Long Range Reconnaissance Imager (LORRI) mounted on NASA’s New Horizons yielded the most accurate measurement of the cosmic optical background (COB). The reported COB intensity is $11.16\pm 1.35$ $\mathrm{nW/m^{2}/sr}$ at a pivot wavelength $\lambda_{piv}=0.608\,\mu\mathrm{m}$ observed in the range $0.4\,\mu\mathrm{m}\lesssim\lambda\lesssim 0.9\,\mu\mathrm{m}$. After subtracting the measured intensity from the deep Hubble Space Telescope count, diffused galactic light, and scattered light from bright star foregrounds, an anomalous intensity of $2.99\pm 2.03\leavevmode\nobreak\ \mathrm{nW/m^{2}/sr}$ has been found. We considered radiatively decaying sterile neutrinos of keV mass scale, as dark matter candidate, that could contribute to this anomalous reported intensity. Using this, we derive upper bounds on the sterile-to-sterile transition magnetic moment. We find that sterile neutrinos with mass of $\mathcal{O}(\rm keV)$ scale take values of the transition magnetic moment in the range $3\times 10^{-13}\,\rm eV^{-1}-10^{-9}\,\rm eV^{-1}$ to explain the anomalous intensity of $2.99\pm 2.03\,\rm nW/m^{2}/sr$.
Sterile-to-sterile transition magnetic moment, Low Energy Effective Field Theory (LEFT)
@@ I Introduction
Line-intensity mapping (LIM) is an observational method for extragalactic astronomy and cosmology that measures the integrated intensity of spectral lines emitted from galaxies and the intergalactic medium at a specific observed frequency Kovetz et al. (2017, 2020). Since LIM experiments capture information from all incoming photons, they have the potential to detect electromagnetic radiation produced by dark matter decays directly. For instance, in Ref. Grin et al. (2007), authors have derived a stringent upper bound on the axion-photon coupling strength by searching for optical line emission produced from decaying axions. In Ref. Gong et al. (2016), authors have shown that photons produced from decaying axions with eV mass can explain the anisotropy of near-IR extragalactic background light. In Ref. Creque-Sarbinowski and Kamionkowski (2018), the authors have proposed that LIM experiments can detect photons produced from decaying and annihilating dark matter particles. However, measuring the extragalactic background light in the optical band is a difficult task. The primary reason is mitigating the overwhelming foreground, which requires accurate modelling and well-calibrated instruments Bernal et al. (2022). In Ref. Zemcov et al. (2017a), authors have suggested observing such radiations in the optical spectrum—the cosmic optical background (COB) is possible if one can eliminate the foreground radiations, primarily solar in nature.
@@
The Long Range Reconnaissance Imager (LORRI) mounted on NASA New Horizons spacecraft, currently operating at a distance $51.3$ astronomical unit (AU) from the Sun, has been measuring the COB photon intensity in the wavelength range $0.4\,\mu\mathrm{m}\lesssim\lambda\lesssim 0.9\,\mu\mathrm{m}$ $(\sim 1.3\leavevmode\nobreak\ \rm eV-3\leavevmode\nobreak\ \,\rm eV)$ Zemcov et al. (2017b); Lauer et al. (2021, 2022); Symons et al. (2023); Postman et al. (2024). In $2021$, LORRI operating at 42-45 AU from the Sun reported an optical intensity from unknown sources of $8.8\pm 4.9\,\rm nW/m^{2}/sr$ - $11.9\pm 4.6\,\rm nW/m^{2}/sr$ Lauer et al. (2021). Later, it yielded an anomalous intensity of COB, $8.06\pm 1.92\,\rm nW/m^{2}/sr$ while operating at $51.3$ AU Lauer et al. (2022). In Ref. Lauer et al. (2022), the authors considered diffused galactic light (DGL) from the Milky Way and scattered light from bright stars (SSL) contributing to the total sky level. Further, they subtracted the intensity from integrated light of external galaxies (IGL), known as deep Hubble galaxy count, $8.17\pm 1.18\,\rm nW/m^{2}/sr$ from the total COB level to report the anomalous intensity Lauer et al. (2022). Later, this result was also supported by an independent research following blind methodology Symons et al. (2023). However, in a recent analysis Postman et al. (2024), the authors have reworked the analysis of the foregrounds from the DGL and found that the contribution of DGL to the COB intensity is greater compared to the one found in the article Lauer et al. (2022). Thus yielding an anomalous intensity $2.99\pm 2.03\,\rm{nW/m^{2}/sr}$ in the optical spectrum, relatively modest to the IGL intensity Postman et al. (2024).
@@
This anomalous photon excess may indicate the presence of physics beyond the Standard Model. Various studies have explored such possibilities, for e.g., photons emitted from decaying axions with masses in the range of $\sim 5\leavevmode\nobreak\ \rm eV-25\leavevmode\nobreak\ \rm eV$ Bernal et al. (2022); Nakayama and Yin (2022); Carenza et al. (2023). Additionally, decaying sterile neutrinos could produce a similar photon signal through their decay into photons and neutrinos.
@@
Short-baseline neutrino experiments have hinted at the existence of additional neutrinos, and numerous studies have investigated sterile neutrinos as a potential dark matter candidate Ho and Scherrer (2013). The investigation of photon radiated from the sterile neutrinos can provide an exciting probe for the beyond standard model containing sterile neutrinos. Motivated by these possibilities, we analyse photon flux originating from decaying sterile neutrinos. We also explore it as the source of the observed anomalous intensity.
@@
In this work, we focus on the keV mass scale of sterile neutrinos, as they are considered a viable dark matter candidate. Rather than considering a specific model, we adopt an effective field theory approach Beltrán et al. (2024), which makes our analysis more general and less dependent on particular model assumptions. We investigate the potential role of the keV-mass scale sterile neutrinos in contributing to the reported anomalous COB intensity, thus deriving upper bounds on the sterile-to-sterile transition magnetic moment by varying the $\Delta m$ values from $2\,\rm eV$ to $10\,\rm eV$. Future experiments like the SPHEREx telescope surveying the sky at wavelengths $0.75-5\,\mu m$ might shed light on the properties of radiatively decaying sterile neutrinos Crill et al. (2020); Doré et al. (2014, 2016, 2018).
@@
The paper is organised as follows. In Section II, we discuss the radiative decay process of sterile neutrinos through the sterile-to-sterile transition magnetic dipole moment in the low-energy effective field theory. In Section III, we derive the mean specific intensity for the keV mass range of decaying sterile neutrinos. In Section IV, we have analyzed our results by determining the minimal required sterile-to-sterile transition magnetic moment and decay width of keV-mass scale sterile neutrinos. In this section, we consider the two COB measurements; one is anomalous COB excess intensity, $8.06\pm 1.92\,\rm nW/m^{2}/sr$ and another one is the rest intensity of COB, $2.99\pm 2.03\,\rm nW/m^{2}/sr$. Here, we also derive the novel constraints on the sterile-to-sterile transition magnetic moment through the contribution of the COB intesnity from the optical to X-rays region. In sectionV, we explain the UV origin of the small mass splitting of sterile neutrinos. In Section VI, we present our conclusions with an outlook for future works.
@@
II Radiative Decay of Sterile Neutrino through Transition Magnetic Moment
@@
We consider Standard Model augmented with sterile neutrinos in the effective field theory approach. Assuming sterile neutrino mass scales well below the electroweak scale, we can write the effective Lagrangian density that contributes to the magnetic moment of the neutrino sector:
$\mathcal{L}_{N_{R}LEFT}\supset d_{\rm NN\gamma}^{ij}\mathcal{O}_{\rm NN\gamma}^{ij}+d_{\nu N\gamma}^{\beta i}\mathcal{O}_{\nu N\gamma}^{\beta i}+\text{h.c.}\,$, (1)
@@
where, $d_{\rm NN\gamma}$ represents the sterile-to-sterile transition magnetic moment and $d_{\rm\nu N\gamma}$ denotes the active-to-sterile transition magnetic moment, respectively. The term $\mathcal{O}_{\rm NN\gamma}^{ij}=(\bar{N}_{Ri}^{c}\sigma_{\mu\nu}N_{Rj})F^{\mu\nu}$, $\mathcal{O}_{\nu N\gamma}^{\alpha i}=(\bar{\nu}_{L\alpha}^{c}\sigma_{\mu\nu}N_{Ri})F^{\mu\nu}$ represent the effective field operators, where $F^{\mu\nu}$ is the electromagnetic field strength tensor. The electric and magnetic field off-diagonal non-zero transition dipole moment components give rise to radiative decay through two different sterile states. Therefore, in this radiative decay process ($m_{1}\rightarrow m_{2}+\gamma$), the decay width of sterile neutrino can be written as Atre et al. (2009); Bondarenko et al. (2018); Balantekin and Vassh (2014)
@@
From Eq. (3), we can derive the sterile-to-sterile transition magnetic moment $d_{\rm NN\gamma}$.
There are several experimental upper bounds on the neutrino magnetic moment. Recent studies of the NOMAND neutrino detector at CERN put constraints on the active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}\lesssim(10^{-6}-10^{-9})\mu_{B}$ Gninenko and Krasnikov (1999). Here, $\mu_{B}\sim 3\times 10^{-7}\,\rm eV^{-1}$ represents the Bohr magneton. In an anti-neutrino electron scattering experiment, by examining the electron recoil spectra, it was found that the active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}\lesssim 10^{-9}\mu_{B}$ Derbin et al. (1993). The Super-Kamiokande experiment put an upper bound limit on active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}\lesssim 1.1\times 10^{-10}\mu_{B}$ Liu et al. (2004), and the Borexino collaboration put the upper bound limit at $90\%$ confidence level as $5.4\times 10^{-11}\mu_{B}$ Agostini et al. (2017); Arpesella et al. (2008). MUNU collaboration has found the upper bound on active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}\lesssim 9.0\times 10^{-11}\mu_{B}$ Babu et al. (2020), whereas TEXONO collaboration has obtained this active-to-sterile transition magnetic moment upper bound limit $d_{\rm\nu N\gamma}\lesssim 7.4\times 10^{-11}\mu_{B}$ Balantekin and Vassh (2014). The Planck+BAO put the constraint on the active-to-sterile transition magnetic moment, and it varies from $3.7\times 10^{-11}\mu_{B}$ to $9.1\times 10^{-12}\mu_{B}$ Carenza et al. (2024). Recent studies analysing the cooling of red giant stars have found an upper limit on active-to-sterile transition magnetic moments $d_{\rm\nu N\gamma}\lesssim 1.2\times 10^{-12}\mu_{B}$ Capozzi and Raffelt (2020). The anomalous stellar cooling due to plasmon decay has been found an upper limit on the active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}$ is $d_{\rm\nu N\gamma}\lesssim 3\times 10^{-12}\mu_{B}$ Alexander (2016).
@@ III Mean specific intensity of decaying sterile neutrinos
where $h$ and $c$ are Planck’s constant and the speed of light in vacuum, respectively. We define the second term on the right-hand side of Eq. (4) as $x^{-1}=\frac{2}{\left(1-m_{\xi}^{2}/m_{\chi}^{2}\right)}$. We note that for a two-photon decay process, $x=1/2$. The wavelength of photon produced from a particle decaying at redshift $z$ will be observed today as $\lambda_{obs}=(1+z)\lambda_{e}$. If the intergalactic gas does not absorb these photons along the line of sight, then they will be observed today as the extragalactic background light. The redshift dependence of the energy density of decaying dark matter particles is given by $\rho_{\chi}(t)=\rho_{\chi}(0)(1+z)^{3}\mathrm{e}^{-\Gamma_{\chi}t}$, where $\Gamma_{\chi}$ represents the decay width, $t$ is the time variable, and $\rho_{\chi}(0)$ is the dark matter energy density today. The specific intensity of the photons $I_{\nu}$ observed at frequency $\nu_{\rm obs}$ today is given by Creque-Sarbinowski and Kamionkowski (2018)
$I_{\nu}=\dfrac{1}{4\pi}\int_{0}^{\infty}{dz}\dfrac{c}{H(z)}\dfrac{\epsilon_{\nu_{\rm obs}}(z)}{(1+z)^{4}},$ (5)
where $\epsilon_{\nu_{\rm obs}}$ is the specific luminosity density, and $H(z)$ is the Hubble parameter. For a Dirac delta decaying profile, $\epsilon_{\nu_{\rm obs}}$ is expressed as Chen and Kamionkowski (2004); Creque-Sarbinowski and Kamionkowski (2018)
@@
From the above equation, we can observe that $I_{\lambda}$ is directly proportional to the decay width and the energy of the released photon. However, it is inversely related to the mass of the decaying particle and decays over redshifts due to the expansion of the universe.